# Can General-Purpose Omnimodels Compete with Specialists? A Case Study in Medical Image Segmentation

### v1

@@ Abstract
The emergence of powerful, general-purpose omnimodels capable of processing diverse data modalities has raised a critical question: can these “jack-of-all-trades” systems perform on par with highly specialized models in knowledge-intensive domains? This work investigates this question within the high-stakes field of medical image segmentation. We conduct a comparative study analyzing the zero-shot performance of a state-of-the-art omnimodel (Gemini 2.5 Pro, the “Nano Banana” model) against domain-specific deep learning models on three distinct tasks: polyp (endoscopy), retinal vessel (fundus), and breast tumor segmentation (ultrasound). Our study focuses on performance at the extremes by curating subsets of the “easiest” and “hardest” cases based on the specialist models’ accuracy. Our findings reveal a nuanced and task-dependent landscape. For polyp and breast tumor segmentation, specialist models excel on easy samples, but the omnimodel demonstrates greater robustness on hard samples where specialists fail catastrophically. Conversely, for the fine-grained task of retinal vessel segmentation, the specialist model maintains superior performance across both easy and hard cases. Intriguingly, qualitative analysis suggests omnimodels may possess higher sensitivity, identifying subtle anatomical features missed by human annotators. Our results indicate that while current omnimodels are not yet a universal replacement for specialists, their unique strengths suggest a potential complementary role with specialist models, particularly in enhancing robustness on challenging edge cases.
@@ 1 Introduction
In this paper, we present three case studies involving polyp segmentation in endoscopic images, retinal vessel segmentation in fundus images, and breast tumor segmentation in ultrasound images, where we compare the zero-shot performance of Gemini 2.5 Pro with specialist models. For tasks involving relatively salient objects, even in noisy modalities like ultrasound, the omnimodel shows surprising robustness on hard cases that cause specialist models to fail completely. In contrast, for the intricate task of retinal vessel segmentation, the specialist model’s architectural priors and focused training give it a decisive advantage across all levels of difficulty. Furthermore, our qualitative analysis uncovers instances where the omnimodel appears to identify fine anatomical structures missed in the ground-truth annotations, suggesting a potential for higher sensitivity. These results provide a nuanced perspective on the current capabilities of omnimodels in medicine, highlighting both their potential and their present limitations.
@@ 3 Experimental Design
  • Omnimodel: As a generalist state-of-the-art model, we adopted Gemini 2.5 Pro, leveraging its “Create images” tool (using the “Nano Banana” model) to assess zero-shot segmentation performance without any task-specific fine-tuning.
@@
Prompting Strategy. We interacted with Gemini 2.5 Pro via its image generation function using task-specific prompts designed to elicit a binary segmentation mask.
@@ 4.1 Polyp Segmentation in Endoscopic Images
Our first case study evaluated performance on the CVC-ColonDB dataset. Table 1 presents the quantitative comparison between the specialist HSNet and the generalist Gemini 2.5 Pro on the stratified easy and hard sample sets.
On the easy samples, the specialist HSNet model demonstrates superior performance, as expected. It achieves near-perfect scores with a Dice of 97.4% and its extremely low HD95 (6.9 pixels) confirms its high boundary precision. While Gemini 2.5 Pro’s performance is lower, its Dice score of 87.7% is still commendable for a zero-shot task, showcasing its strong general-purpose capabilities. Visualizations in Figure 1 corroborate these findings, showing HSNet’s masks closely aligning with the ground truth.
@@
Omnimodel (Gemini 2.5 Pro) 87.7 87.4
@@
Omnimodel (Gemini 2.5 Pro) 23.6 304.6
The performance dynamic reverses on the hard samples. HSNet’s performance collapses, with its Dice score plummeting to a mere 4.3%, indicating a complete failure to segment the polyp, as visualized in Figure 2. In contrast, Gemini 2.5 Pro achieves a Dice score of 23.6%. While modest, this result is substantially better than HSNet’s and suggests that the omnimodel possesses greater robustness and generalization when faced with ambiguous or out-of-distribution examples that cause the specialist to fail. This highlights a potential advantage of omnimodels in handling challenging edge cases.
@@ 4.2 Retinal Vessel Segmentation in Fundus Images
On both easy and hard samples, the specialized U-Net model maintains a decisive performance advantage. For easy cases, the U-Net achieves a Dice score of 91.5%, significantly outperforming Gemini 2.5 Pro’s 62.4%. Unlike the polyp task, this performance gap persists on the hard samples. The U-Net achieves a Dice score of 58.7% on these challenging images, more than double Gemini 2.5 Pro’s score of 28.3%.
@@
Omnimodel (Gemini 2.5 Pro) 62.4 111.0
@@
Omnimodel (Gemini 2.5 Pro) 28.3 408.9
@@
Despite its lower quantitative scores, a notable qualitative finding emerged. As shown in Figure 5, Gemini 2.5 Pro often identifies fine vessels that are absent in the ground truth but become visible in contrast-enhanced versions of the original images. This suggests the omnimodel may possess a higher sensitivity for detecting minute details overlooked during manual annotation, a potentially valuable characteristic.
@@
Figure 5: Zoomed-in qualitative comparison of retinal vessel segmentation. Gemini 2.5 Pro often identifies fine vessels (third column) missed in the ground truth (second column) but visible with contrast enhancement (fourth column).
@@ 4.3 Breast Tumor Segmentation in Ultrasound Images
On easy samples, the specialist Mask2Former model demonstrates clear superiority, achieving a high Dice score of 95.2% compared to Gemini 2.5 Pro’s 70.5%. This confirms the specialist’s effectiveness on straightforward, in-distribution examples.
However, on the hard samples, we again observe a role reversal. Mask2Former’s performance catastrophically fails, dropping to a Dice score of just 3.1%. This indicates its inability to handle challenging cases with ambiguous boundaries or unusual presentations. In contrast, Gemini 2.5 Pro shows better resilience, maintaining a Dice score of 21.1%. Although this is not a high score, it reinforces the observation that the omnimodel’s broad knowledge base provides a degree of robustness against the types of edge cases that cause narrowly trained models to fail. This result from a challenging, noisy modality further strengthens the hypothesis that omnimodels may have a unique advantage in handling difficult, out-of-distribution samples.
@@
Omnimodel (Gemini 2.5 Pro) 70.5 121.3
@@
Omnimodel (Gemini 2.5 Pro) 21.1 198.1

### v2

@@ Abstract
The emergence of powerful, general-purpose omnimodels capable of processing diverse data modalities has raised a critical question: can these “jack-of-all-trades” systems perform on par with highly specialized models in knowledge-intensive domains? This work investigates this question within the high-stakes field of medical image segmentation. We conduct a comparative study analyzing the zero-shot performance of a state-of-the-art omnimodel (Gemini, the “Nano Banana” model) against domain-specific deep learning models on three distinct tasks: polyp (endoscopy), retinal vessel (fundus), and breast tumor segmentation (ultrasound). Our study focuses on performance at the extremes by curating subsets of the “easiest” and “hardest” cases based on the specialist models’ accuracy. Our findings reveal a nuanced and task-dependent landscape. For polyp and breast tumor segmentation, specialist models excel on easy samples, but the omnimodel demonstrates greater robustness on hard samples where specialists fail catastrophically. Conversely, for the fine-grained task of retinal vessel segmentation, the specialist model maintains superior performance across both easy and hard cases. Intriguingly, qualitative analysis suggests omnimodels may possess higher sensitivity, identifying subtle anatomical features missed by human annotators. Our results indicate that while current omnimodels are not yet a universal replacement for specialists, their unique strengths suggest a potential complementary role with specialist models, particularly in enhancing robustness on challenging edge cases.
@@ 1 Introduction
In this paper, we present three case studies involving polyp segmentation in endoscopic images, retinal vessel segmentation in fundus images, and breast tumor segmentation in ultrasound images, where we compare the zero-shot performance of Gemini with specialist models. For tasks involving relatively salient objects, even in noisy modalities like ultrasound, the omnimodel shows surprising robustness on hard cases that cause specialist models to fail completely. In contrast, for the intricate task of retinal vessel segmentation, the specialist model’s architectural priors and focused training give it a decisive advantage across all levels of difficulty. Furthermore, our qualitative analysis uncovers instances where the omnimodel appears to identify fine anatomical structures missed in the ground-truth annotations, suggesting a potential for higher sensitivity. These results provide a nuanced perspective on the current capabilities of omnimodels in medicine, highlighting both their potential and their present limitations.
@@ 3 Experimental Design
  • Omnimodel: As a generalist state-of-the-art model, we adopted Gemini, using its “Create images” tool (using the “Nano Banana” model) to assess zero-shot segmentation performance without any task-specific fine-tuning.
@@
Prompting Strategy. We interacted with Gemini via its image generation function using task-specific prompts designed to elicit a binary segmentation mask.
@@ 4.1 Polyp Segmentation in Endoscopic Images
Our first case study evaluated performance on the CVC-ColonDB dataset. Table 1 presents the quantitative comparison between the specialist HSNet and the generalist Gemini on the stratified easy and hard sample sets.
On the easy samples, the specialist HSNet model demonstrates superior performance, as expected. It achieves near-perfect scores with a Dice of 97.4% and its extremely low HD95 (6.9 pixels) confirms its high boundary precision. While Gemini Nano Banana’s performance is lower, its Dice score of 87.7% is still commendable for a zero-shot task, showcasing its strong general-purpose capabilities. Visualizations in Figure 1 corroborate these findings, showing HSNet’s masks closely aligning with the ground truth.
@@
Omnimodel (Gemini Nano Banana) 87.7 87.4
@@
Omnimodel (Gemini Nano Banana) 23.6 304.6
The performance dynamic reverses on the hard samples. HSNet’s performance collapses, with its Dice score plummeting to a mere 4.3%, indicating a complete failure to segment the polyp, as visualized in Figure 2. In contrast, Gemini Nano Banana achieves a Dice score of 23.6%. While modest, this result is substantially better than HSNet’s and suggests that the omnimodel possesses greater robustness and generalization when faced with ambiguous or out-of-distribution examples that cause the specialist to fail. This highlights a potential advantage of omnimodels in handling challenging edge cases.
@@ 4.2 Retinal Vessel Segmentation in Fundus Images
On both easy and hard samples, the specialized U-Net model maintains a decisive performance advantage. For easy cases, the U-Net achieves a Dice score of 91.5%, significantly outperforming Gemini Nano Banana’s 62.4%. Unlike the polyp task, this performance gap persists on the hard samples. The U-Net achieves a Dice score of 58.7% on these challenging images, more than double Gemini Nano Banana’s score of 28.3%.
@@
Omnimodel (Gemini Nano Banana) 62.4 111.0
@@
Omnimodel (Gemini Nano Banana) 28.3 408.9
@@
Despite its lower quantitative scores, a notable qualitative finding emerged. As shown in Figure 5, Gemini Nano Banana often identifies fine vessels that are absent in the ground truth but become visible in contrast-enhanced versions of the original images. This suggests the omnimodel may possess a higher sensitivity for detecting minute details overlooked during manual annotation, a potentially valuable characteristic.
@@
Figure 5: Zoomed-in qualitative comparison of retinal vessel segmentation. Gemini Nano Banana often identifies fine vessels (third column) missed in the ground truth (second column) but visible with contrast enhancement (fourth column).
@@ 4.3 Breast Tumor Segmentation in Ultrasound Images
On easy samples, the specialist Mask2Former model demonstrates clear superiority, achieving a high Dice score of 95.2% compared to Gemini Nano Banana’s 70.5%. This confirms the specialist’s effectiveness on straightforward, in-distribution examples.
However, on the hard samples, we again observe a role reversal. Mask2Former’s performance catastrophically fails, dropping to a Dice score of just 3.1%. This indicates its inability to handle challenging cases with ambiguous boundaries or unusual presentations. In contrast, Gemini Nano Banana shows better resilience, maintaining a Dice score of 21.1%. Although this is not a high score, it reinforces the observation that the omnimodel’s broad knowledge base provides a degree of robustness against the types of edge cases that cause narrowly trained models to fail. This result from a challenging, noisy modality further strengthens the hypothesis that omnimodels may have a unique advantage in handling difficult, out-of-distribution samples.
@@
Omnimodel (Gemini Nano Banana) 70.5 121.3
@@
Omnimodel (Gemini Nano Banana) 21.1 198.1